### v1

@@ 2.4 Asymptotic solution: Compressible neo–Hookean bilayer system
$+d_{3}r+d_{4}r^{4/3}+d_{5}r^{5/3}+\pazocal{O}(r^{2}),$ (4)
where the coefficients $c_{4},\cdots,c_{6},d_{3},\cdots,d_{5}$ are shown in Appendix B: Coefficients in asymptotic expressions.
[…]
In Fig. 6(a-b), we plot the variations of $\lambda_{c}$ and $k_{c}h$ with the Poisson ratio of the film ($0.450<\nu_{f}<$ 0.45) for different values of shear moduli contrast $\mu_{f}/\mu_{s}$ using Eq. 3-Eq. 4, because they give curves instantly compared to long computations.
Figure 6: Using the asymptotic expressions Eq. 3-Eq. 4 for compressible neo-Hookean materials ($\alpha_{s}=\alpha_{f}=1$ in Eq. 1). The substrate is quasi-incompressible ($\nu_{s}=0.495$). (a-b): Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}h$ with the Poisson ratio of film layer ($k_{c}$ is the critical wavenumber and $h$ is the deformed film thickness) and for contrast in shear moduli between the layers $(\nicefrac{{\mu_{f}}}{{\mu_{s}}}=\{50,100,500,1000\})$.
3 Methods
3.1 Semi-analytical treatment
We use Eq. 1 to determine the principal components of the Cauchy stress $\mathbf{T}=J^{-1}\mathbf{F}\left(\partial W/\partial\mathbf{F}\right)$ as
$T_{i}=E\frac{J^{-(2\alpha+1)}\left(-\alpha\lambda_{i}^{2}+(\alpha-1+\alpha\lambda_{i}^{4})J^{2\alpha}-(\alpha-1)\lambda_{i}^{2}J^{4\alpha}\right)}{2\lambda_{i}^{2}\left(1+\nu\right)},$ (5)
where the $\lambda_{i}$ are the principal stretch ratios and $J=\det\mathbf{F}$.
The substrate is under uniaxial tension along X, so that $T^{s}_{22}=T^{s}_{33}=0$, which gives the following deformation gradient,
$\mathbf{F}^{{s}}=\text{diag}\left(\lambda_{1},\lambda_{1}^{-\nu_{{s}}},\lambda_{1}^{-\nu_{{s}}}\right).$ (6)
The film and substrate are perfectly bonded, so that $\lambda_{1}^{f}=\lambda_{1}^{s}$, $\lambda_{3}^{f}=\lambda_{3}^{s}$, and $T^{f}_{22}=T^{s}_{22}=0$. We then find that
$\mathbf{F}^{{f}}=\text{diag}\left(\lambda_{1},\lambda_{1}^{\frac{\nu_{{f}}(\nu_{{s}}-1)}{(1-\nu_{{f}})}},\lambda_{1}^{-\nu_{{s}}}\right).$ (7)
To find the critical state of buckling, a small-amplitude mechanical displacement $\mathbf{u}$ is superimposed on the finite deformations. The incremental equations of equilibrium read
$\pazocal{A}_{0jilm}^{{s}}u_{m,lj}^{{s}}=0,$ $-\infty<y<0,$
$\pazocal{A}_{0jilm}^{{f}}u_{m,lj}^{{f}}=0,$ $0<y<h,$ (8)
where the commas denote differentiation with respect to the coordinates, $h$ is the current thickness of the film, and $\pazocal{A}_{0}$ is the fourth-order tensor of the instantaneous elastic moduli, with components
$\pazocal{A}_{0jilm}=J^{-1}F_{j\alpha}\frac{\partial^{2}W}{\partial F_{i\alpha}\partial F_{m\beta}}F_{l\beta}.$ (9)
We look for solutions of the forms:
$u_{z}=e^{sy}\sin(kz),\qquad u_{y}=e^{sy}\cos(kz),$ (10)
where $s$ is the attenuation coefficient and $k$ is the wavenumber of the sinusoidal wrinkles. Substitution into Eq. 8 leads to an eigenproblem, with characteristic equation a bicubic in $s$. In the film, the general solution is of the form
$\begin{split}u_{z}^{{f}}&=\Big{(}\sum_{i=1}^{4}\pazocal{V}_{i}e^{s_{i}y}\Big{)}\sin(kz),\quad u_{y}^{{f}}=\Big{(}\sum_{i=1}^{4}\pazocal{V}_{i}e^{s_{i}y}\Big{)}\cos(kz),\end{split}$ (11)
where $s_{1},\cdots,s_{4}$ are the eigenvalues, and $\pazocal{V}_{1},\cdots,\pazocal{V}_{4}$ are constants. In the substrate, the stretch ratios along the Y- and Z- directions are equal, and $s=1$ is a repeated eigenvalue. The other repeated root, $s=-1$ is discarded to enforce decay. There, the solution is thus of the form
$\begin{split}u_{z}^{{s}}&=(\pazocal{U}_{1}+\pazocal{U}_{2}y)e^{y}\sin(kz),\qquad u_{y}^{{s}}=(\pazocal{U}_{1}+\pazocal{U}_{2}y)e^{y}\cos(kz),\end{split}$ (12)
where $\pazocal{U}_{1},\pazocal{U}_{2}$ are constants.
By applying the traction-free boundary conditions
$\pazocal{A}_{02ilm}^{{f}}u_{m,l}^{{f}}=0,\qquad y=h,$ (13a)
and the continuity conditions
$\pazocal{A}_{02ilm}^{{f}}u_{m,l}^{{f}}=\pazocal{A}_{02ilm}^{{s}}u_{m,l}^{{s}},\qquad u_{i}^{{f}}=u_{i}^{{s}},\qquad y=0,$ (13b)
we obtain six homogeneous equations for $\{{\pazocal{U}_{1},\pazocal{U}_{2},\pazocal{V}_{1},\cdots,\pazocal{V}_{4}\}}$. The bifurcation condition is then given by equating the determinant of a $6\times 6$ coefficient matrix to 0.
Figure 7: (a) Undeformed and (b) deformed configurations of auxetic stiff film with orthogonal-oval shaped patterns bonded to a compliant substrate. The base materials for film and substrate are incompressible. Periodic boundary conditions are applied on the extreme faces of the domain along X and Z and the system is subjected to uni-axial tension along X. In (b), we observe that the wrinkles are generated parallel to the direction of tension. The critical strain for buckling is $\varepsilon_{c}\simeq 0.274.$
3.2 Finite Element simulations: 3D models
We first checked the accuracy of our periodic boundary conditions (PBCs) code (Python script) to reproduce known results for wrinkling under plane-strain compression. Hence, we performed a linear buckling analysis in ABAQUS on an incompressible neo-Hookean bilayer periodic along the X-direction, see Appendix A: Validation of PBCs code – 2D incompressible neo-Hookean bilayer model, and recovered the results of [42].
We then perform a series of finite-element simulations for 3D Blatz-Ko model bilayers under uni-axial tension. The dimensions (X, Y, Z) of the bilayer structure are $15\times 45\times 5.684$ units with ${h_{{s}}}/{h_{{f}}}=299$ when the substrate is incompressible, and $15\times 80\times 4.504$ units with ${h_{{s}}}/{h_{{f}}}=399$ when the film is highly auxetic.
The dimensions along the Z-direction are integer multiples of the wavelength (calculated using semi-analytical results from Mathematica) when $(\nu_{f},\nu_{s})=\{0.3$, 0.495}0.495\}0.495 } and $\{-0.95,-0.8\}$. For the other $(\nu_{f},\nu_{s})$ pairs, we noticed that small variations in the depth of the structure did not notably alter the values of critical stretch (changes in the third decimal point). To save computational time (without compromising the efficiency of the solutions), we kept the dimensions $15\times 45\times 5.684$ units and $15\times 80\times 4.504$ units when the substrate is incompressible (Section 2.2) and when the film is highly auxetic (Section 2.3).
For the neo-Hookean bilayers ($\alpha_{f}=\alpha_{s}=1$), the dimensions are $6\times 45\times 1.85$ units with ${h_{{s}}}/{h_{{f}}}=299$.
We apply PBCs on the extreme faces in the X- and Z-directions of the domain using a user-defined Python script file. Both the film and substrate are modelled using a UHYPER subroutine with the Blatz-Ko strain energy function. We use a 20-node brick (hexahedral) element with quadratic interpolation and reduced integration (C3D20R) for both film and substrate; for the limiting case of incompressibility, we used a C3D20RH element. For the Blatz-Ko bilayers, 34,200 and 17,130 mesh elements are used for the cases of an incompressible substrate and a highly auxetic film, respectively. For the neo-Hookean bilayer system, 3,240 mesh elements are used.
We find that the numerical results obtained using different approaches in ABAQUS and Mathematica software match well for a wide range of material parameters, see Fig. 4. When $\alpha_{f}=\alpha_{s}=1$ (compressible neo-Hookean bilayers), the results using asymptotic expressions agree well with the numerical simulations, see Fig. 5.
4 Conclusions
We investigated the possibility of harnessing wrinkles parallel to the direction of applied tension in 3D isotropic compressible bilayers subject to large elongations. We paid particular attention to the cases when the substrate is quasi-incompressible and when the film is highly auxetic.
We used a semi-analytical approach with Mathematica to predict the onset of wrinkling, and user-defined Python scripts to apply periodic boundary conditions and UHYPER subroutine for Blatz-Ko material models to perform linear buckling analysis in Finite Element simulations with ABAQUS.
For compressible neo–Hookean bilayer systems, we derived asymptotic expressions for critical stretch ratio and critical wavenumber, which can be used under finite strains to determine the Young modulus of the film layer for buckling-based metrology applications.
We found that wrinkles can be obtained only when the Poisson ratio of the substrate is greater than that of the film. When the Poisson ratios of film and substrate converge to a common value, the critical stretch ratio shoots up sharply and the wavelength of wrinkles is high. In the limit, the wrinkles are not present because there are no compressive stresses developing when the lateral expansions are the same for the film and substrate. Through multiple simulations we showed that by varying the material properties, we can harness or delay the onset of wrinkles.
@@ 4 Conclusions
Some of the limitations of our work include the consideration of isotropic strain energy function for auxetic materials, and deformation-independent material properties. Functional-grading of auxetics could also be explored with the methods presented in this study, see some preliminary works on harnessing instabilities in functionally-graded auxetic materials using tension-field theory [43, 44] and their applications [45, 46, 47].
Another area of interest concerns instabilities in materials with auxetic patterns. Indeed, auxetic properties can be obtained at a continuum level with a careful design of holes or voids at the micro-scale [6, 48]. For example, in Fig. 7 we performed linear buckling analysis when the substrate is covered with an initially incompressible film made auxetic by orthogonal oval-shaped voids, leading to an effective Poisson ratio ranging between -0.2 and 0 [9]. We see that the critical stretch ratio $\lambda_{c}\simeq 1.274$ matches well (within 5%) with the results in Fig. 4.
@@
This paper’s analysis could play a critical role in manufacturing and testing the applicability of auxetic hydrogel organ patches [9] and skin grafts [49].
Appendix A: Validation of PBCs code – 2D incompressible neo-Hookean bilayer model
Figure A.1: (a) Schematic representation of an incompressible neo-Hookean stiff film/soft substrate bilayer system under uni-axial compression. On the left edge of the domain, displacement and shear traction are set at zero. On the bottom edge of the domain, roller support restricts vertical displacement and shear traction, and the top surface of the film is traction-free. (b) Linear buckling solutions with critical strains $\varepsilon_{c}$ when ${\mu_{f}}/{\mu_{s}}=30,1000$, in line with the predictions of Cao and Hutchinson [42].
Under uni-axial, plane-strain compression, the theoretical critical strain is $\varepsilon_{c}=\nicefrac{{1}}{{4}}\left({3\mu_{s}}/{\mu_{f}}\right)^{\nicefrac{{2}}{{3}}}$ [42].
In Fig. A.1, we applied PBCs on the left and right edges of the domain, roller support on the bottom edge of the substrate, perfect bonding between film and substrate layers, and traction-free condition on the top surface of the film in the two cases ${\mu_{f}}/{\mu_{s}}=30,1000$. We find $\varepsilon_{c}=0.053,0.0052$, respectively, matching well with the theoretical solutions.
Having validated our 1D PBCs script along the X-direction on a 2D model, we then extended the PBCs to two dimensions, along the X- and Z-directions, to perform a linear buckling analysis on 3D bilayer systems under uniaxial tension.
The height ratio of substrate to film layers was taken as ${h_{{s}}}/{h_{{f}}}=163$, with the width of layers being 30.457 units. We modelled both film and substrate with the incompressible neo-Hookean model.
We used a hybrid 8-node plane strain quadrilateral element with quadratic interpolation and reduced integration (CPE8RH) for both film and substrate layers. We also applied periodic boundary conditions on the left and right edges of the domain. The minimum size of the mesh element is lower than the height of the film. We took 12,800 elements, and the numerical results converged and were consistent with the theoretical solutions.
Appendix B: Coefficients in asymptotic expressions
$\begin{split}c_{4}&=\frac{\pazocal{C}\left(45+\nu_{s}c_{4a}+\nu_{f}c_{4b}+\nu_{f}^{2}c_{4c}\right)}{20\times 2^{1/3}\times 3^{2/3}(\nu_{f}-\nu_{s})^{2}(-1+\nu_{s})(-3+4\nu_{s})^{4/3}},\\ c_{5}&=\frac{\pazocal{C}^{2}(-1+\nu_{f})(-1+2\nu_{s})(-3-4\nu_{f}+4\nu_{s})}{2^{2/3}\times 3^{1/3}(\nu_{f}-\nu_{s})^{2}(-3+4\nu_{s})^{5/3}},\\ c_{6}&=\frac{4725-2\nu_{f}^{4}c_{6a}-\nu_{s}c_{6b}+\nu_{f}^{3}c_{6c}+\nu_{f}^{2}c_{6d}+2\nu_{f}c_{6e}}{12600(3-4\nu_{s})^{2}(-1+\nu_{s})^{2}(-\nu_{f}+\nu_{s})^{3}},\\ \pazocal{C}&=\left((1-\nu_{f})(\nu_{s}-1)\right)^{1/3},\end{split}$ (B.1a)
and
$\begin{split}c_{4a}&=-64+(13-4\nu_{s})\nu_{s},\\ c_{4b}&=-71+\nu_{s}(66+(51-26\nu_{s})\nu_{s}),\\ c_{4c}&=56-92\nu_{s}+26\nu_{s}^{2},\\ c_{6a}&=33727+2\nu_{s}(-73474+\nu_{s}(122601\\ &+\nu_{s}(-92584+26681\nu_{s}))),\\ c_{6b}&=4410+\nu_{s}(82574+\nu_{s}(-288226\\ &+\nu_{s}(399999-260716\nu_{s}+67034\nu_{s}^{2}))),\\ c_{6c}&=95848+2\nu_{s}(-138872+\nu_{s}(46253\\ &+\nu_{s}(235988+\nu_{s}(-297317+106724\nu_{s})))),\\ c_{6d}&=-3509+\nu_{s}(-202720+\nu_{s}(841376+\nu_{s}(-1218634\\ &+\nu_{s}(643241+2(22435-53362\nu_{s})\nu_{s})))),\\ c_{6e}&=-11970+\nu_{s}(78479+\nu_{s}(-137927+\nu_{s}(-12652\\ &+\nu_{s}(271583+\nu_{s}(-276527+89714\nu_{s}))))).\end{split}$ (B.1b)
Similarly, the coefficients of higher-order terms in the asymptotic expression for critical wavenumber (Eq. 4) are
$\begin{split}d_{3}&=\frac{4+2\nu_{s}-11\nu_{s}^{2}+\nu_{f}\left(11-32\nu_{s}+26\nu_{s}^{2}\right)}{15\left(3-7\nu_{s}+4\nu_{s}^{2}\right)},\\ d_{4}&=\frac{\left(\frac{2}{3}\right)^{2/3}\pazocal{C}(1+2\nu_{f})(-1+2\nu_{s})}{(-3+4\nu_{s})^{4/3}},\\ d_{5}&=\frac{\pazocal{C}^{2}\left(-6761+\nu_{s}d_{5a}+4\nu_{f}^{2}d_{5b}-2\nu_{f}d_{5c}\right)}{3150\times 2^{2/3}\times 3^{1/3}(-1+\nu_{f})(-1+\nu_{s})^{3}(-3+4\nu_{s})^{5/3}},\end{split}$ (B.2a)
and
$\begin{split}d_{5a}&=21724+\nu_{s}(-19146+\nu_{s}(-2636+7519\nu_{s})),\\ d_{5b}&=646+\nu_{s}(-4964+\nu_{s}(12906\\ &+\nu_{s}(-14204+5791\nu_{s}))),\\ d_{5c}&=-3821+\nu_{s}(7864+\nu_{s}(5844\\ &+\nu_{s}(-22796+13609\nu_{s}))).\end{split}$ (B.2b)
[…]
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
Supplementary data
There is no Supplementary material attached to this article.

### v2

@@ Abstract
$+d_{3}r+d_{4}r^{4/3}+d_{5}r^{5/3}+\mathcal{O}(r^{2}),$ (4)
where the coefficients $c_{4},\cdots,c_{6},d_{3},\cdots,d_{5}$ are shown in the section Asymptotic Analysis in the Supplementary Material. We find that in small deformations our predictions recover the linear-elastic behavior, as described in Equation (2). However, as $\nu_{f}\!\to\!\nu_{s}$, the critical strain becomes large and the linear-elastic results greatly underestimate the numerical results, while our asymptotic expressions in Equations (3) and (4) give good agreement. Figure 4(a) and (b) show how $\lambda_{c}$ and $k_{c}h$ vary with $\nu_{f}$ when $\nu_{s}\!=\!0.495$ (quasi-incompressible substrate). In Figure S.2 (Supplementary Material), we plot the variations of $\lambda_{c}$ and $k_{c}h$ with the Poisson ratio of the film ($0.450<\nu_{f}<$ 0.45) for different values of shear moduli contrast $\mu_{f}/\mu_{s}$ using Equations 3-4.
We now demonstrate how our results can be used to predict wrinkling instabilities in materials with auxetic patterns. Indeed, auxetic properties can be obtained at a continuum level with a careful design of holes or voids at the micro-scale [6, 42]. Figure 5(a) illustrates how the geometric parameters of the orthogonal oval and rectangular voids generate the effective Poisson’s ratios ($\nu_{f}^{\text{eff}}$). Using the inverse design approach outlined by Danesh et al. [43], we generate the geometries for a set of auxetic structures with the desired effective Poisson’s ratios ($\nu_{f}^{\text{eff}}$) in the small strain regime, as depicted in Figure 5(b). We validate the effective properties of these resulting geometries with the Finite Element code, FEAP [44], demonstrating a strong agreement between predictions and desired results (see Supplementary Material: Inverse Design and Homogenization for more details).
Finally, in Figure 5(b), we apply our new numerical method to predict wrinkling in bilayers with different microstructural patterns. For this analysis, we design film layers with orthogonal oval, rectangular, and sinusoidal voids, each tailored to yield specific Poisson’s ratios shown in Figure 5(b). The critical stretch ratio for buckling in bilayers with microstructural patterns aligns well with the homogenized model outcomes presented in Figure 3(d). Indeed, the relative error for the critical stretch between microstructural and the homogenized models is within 6.3%.
In conclusion, in this letter we investigated the possibility of harnessing wrinkles parallel to the direction of applied tension in 3D isotropic compressible bilayers subject to large elongations. We paid particular attention to the cases when the substrate is quasi-incompressible and when the film is highly auxetic. We used a semi-analytical approach to predict the onset of wrinkling. We wrote a new UHYPER subroutine for Blatz–Ko material models and new user-defined Python scripts to simulate wrinkling in compressible bilayers with periodic boundary conditions. These scripts can be used to perform a linear buckling analysis in Finite Element simulations with ABAQUS. For compressible neo–Hookean bilayer systems, we derived asymptotic expressions for critical stretch ratio and critical wavenumber, which can be used under finite strains to determine the Young modulus of the film layer for buckling-based metrology applications. We found that wrinkles can be obtained only when the Poisson ratio of the substrate is greater than that of the film. When the Poisson ratios of film and substrate converge to a common value, the critical stretch ratio shoots up sharply and the wavelength of wrinkles is high. In the limit, the wrinkles are not present because there are no compressive stresses developing when the lateral expansions are the same for the film and substrate. Through multiple simulations we showed that by varying the material properties, we can harness or delay the onset of wrinkles.
Additionally, using inverse analysis, we designed film microstructures to achieve desired effective Poisson’s ratios, validated with Finite Element code, FEAP. The critical stretch ratio for buckling in auxetic structures with microstructural patterns closely aligns with predictions from homogenized models.
@@
Some of the limitations of our work include the consideration of isotropic strain energy function for auxetic structures, and deformation-independent effective material properties. Functional-grading of auxetics could also be explored with the methods presented in this study, see some preliminary works on harnessing instabilities in functionally-graded auxetic materials using tension-field theory [45, 46] and their applications [47, 48, 49].
@@
Ultimately, the method developed in this letter could play a critical role in the manufacturing and testing of auxetic hydrogel organ patches [9] and skin grafts [50]. This approach could specifically guide the design of graft microstructures to prevent undesired wrinkling.